# ORCID: 0009-0007-8992-9682

### v2

@@ 5 Experiment Results
Table 13: Performance of code generation models against docstring perturbations (HumanEval-X and worst case passes)
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS Nominal RP5@1 0.02 0.1 0.09 0.11 0.12 0.13 0.13 0.13 0.13 0.16 0.15 0.16 EnglishInflectionalVariation RP5@1 0.0 0.04 0.03 0.04 0.05 0.06 0.07 0.06 0.05 0.07 0.08 0.05 RD5@1 1.0 0.59 0.62 0.67 0.55 0.52 0.48 0.5 0.62 0.58 0.43 0.7 RR5@1 0.02 0.07 0.05 0.07 0.07 0.07 0.06 0.06 0.08 0.1 0.06 0.12 TenseTransformationFuture RP5@1 0.01 0.05 0.04 0.04 0.07 0.07 0.09 0.08 0.07 0.05 0.08 0.06 RD5@1 0.75 0.47 0.46 0.61 0.45 0.48 0.33 0.35 0.43 0.65 0.43 0.63 RR5@1 0.02 0.07 0.04 0.08 0.05 0.07 0.05 0.05 0.05 0.1 0.06 0.1 ButterFingersPerturbation RP5@1 0.0 0.04 0.04 0.03 0.06 0.05 0.07 0.06 0.04 0.04 0.07 0.06 RD5@1 1.0 0.59 0.46 0.72 0.5 0.57 0.48 0.5 0.67 0.73 0.48 0.63 RR5@1 0.02 0.09 0.05 0.09 0.06 0.07 0.06 0.07 0.09 0.12 0.07 0.1 ChangeCharCase RP5@1 0.0 0.01 0.01 0.02 0.06 0.04 0.05 0.07 0.03 0.04 0.07 0.04 RD5@1 1.0 0.88 0.85 0.78 0.5 0.71 0.62 0.45 0.76 0.73 0.48 0.74 RR5@1 0.02 0.09 0.08 0.1 0.07 0.09 0.08 0.05 0.1 0.12 0.07 0.12 BackTranslation RP5@1 0.0 0.04 0.04 0.04 0.08 0.07 0.09 0.09 0.08 0.06 0.09 0.07 RD5@1 1.0 0.59 0.54 0.61 0.35 0.43 0.33 0.3 0.38 0.62 0.39 0.56 RR5@1 0.02 0.09 0.04 0.08 0.05 0.08 0.07 0.04 0.06 0.1 0.07 0.1 SwapCharactersPerturbation RP5@1 0.0 0.04 0.02 0.02 0.07 0.05 0.09 0.07 0.05 0.04 0.07 0.07 RD5@1 1.0 0.59 0.77 0.78 0.45 0.57 0.33 0.45 0.62 0.73 0.48 0.59 RR5@1 0.02 0.09 0.06 0.09 0.05 0.09 0.05 0.05 0.08 0.12 0.07 0.11 TenseTransformationPast RP5@1 0.0 0.06 0.04 0.04 0.06 0.06 0.07 0.09 0.06 0.04 0.09 0.07 RD5@1 1.0 0.41 0.46 0.61 0.5 0.52 0.43 0.3 0.52 0.77 0.39 0.59 RR5@1 0.02 0.07 0.04 0.08 0.06 0.07 0.05 0.05 0.07 0.12 0.05 0.11 SynonymSubstitution RP5@1 0.01 0.04 0.02 0.04 0.05 0.04 0.04 0.05 0.05 0.04 0.08 0.07 RD5@1 0.5 0.59 0.77 0.61 0.55 0.67 0.67 0.55 0.62 0.73 0.43 0.56 RR5@1 0.02 0.07 0.06 0.09 0.07 0.09 0.09 0.07 0.08 0.12 0.07 0.1 WhitespacePerturbation RP5@1 0.0 0.05 0.02 0.04 0.05 0.04 0.06 0.06 0.05 0.04 0.07 0.05 RD5@1 1.0 0.53 0.69 0.67 0.55 0.67 0.52 0.5 0.62 0.77 0.52 0.67 RR5@1 0.02 0.08 0.05 0.09 0.07 0.1 0.07 0.06 0.08 0.12 0.07 0.12 SynonymInsertion RP5@1 0.01 0.05 0.02 0.03 0.06 0.05 0.09 0.07 0.05 0.07 0.09 0.03 RD5@1 0.75 0.47 0.69 0.72 0.5 0.62 0.33 0.4 0.62 0.54 0.39 0.81 RR5@1 0.02 0.06 0.07 0.09 0.06 0.08 0.05 0.05 0.08 0.1 0.07 0.15
Table 14: Performance of code generation models against function name perturbations. (HumanEval-X and worst case passes)
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi
Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS
Nominal RP5@1 0.02 0.1 0.09 0.11 0.12 0.13 0.13 0.13 0.13 0.16 0.15 0.16
FuncRenameSynonymSub RP5@1 0.01 0.03 0.04 0.04 0.07 0.07 0.08 0.07 0.04 0.05 0.07 0.09
RD5@1 0.75 0.71 0.54 0.67 0.45 0.48 0.38 0.45 0.67 0.69 0.48 0.48
RR5@1 0.02 0.09 0.05 0.07 0.07 0.06 0.05 0.07 0.09 0.11 0.07 0.09
FuncRenameInflectionalVariation RP5@1 0.0 0.03 0.04 0.04 0.07 0.07 0.09 0.07 0.07 0.03 0.09 0.07
RD5@1 1.0 0.71 0.54 0.67 0.4 0.43 0.29 0.45 0.43 0.81 0.39 0.59
RR5@1 0.02 0.07 0.04 0.07 0.06 0.08 0.05 0.05 0.07 0.13 0.05 0.1
FuncRenameChangeChar RP5@1 0.0 0.02 0.05 0.05 0.07 0.06 0.06 0.04 0.07 0.05 0.07 0.07
RD5@1 1.0 0.82 0.31 0.56 0.45 0.52 0.52 0.65 0.43 0.69 0.48 0.56
RR5@1 0.02 0.1 0.05 0.07 0.05 0.08 0.07 0.08 0.07 0.13 0.08 0.12
FuncRenameButterFinger RP5@1 0.0 0.05 0.03 0.03 0.04 0.07 0.07 0.08 0.07 0.04 0.07 0.07
RD5@1 1.0 0.47 0.62 0.72 0.65 0.48 0.43 0.35 0.48 0.73 0.48 0.56
RR5@1 0.02 0.09 0.05 0.08 0.08 0.07 0.05 0.04 0.07 0.12 0.07 0.1
FuncRenameSwapChar RP5@1 0.01 0.05 0.04 0.05 0.06 0.07 0.07 0.07 0.07 0.05 0.09 0.07
RD5@1 0.75 0.53 0.46 0.56 0.5 0.48 0.43 0.4 0.43 0.69 0.35 0.56
RR5@1 0.02 0.07 0.05 0.07 0.06 0.07 0.05 0.06 0.07 0.11 0.06 0.12
FuncRenameSnake/CamelCase RP5@1 0.01 0.06 0.04 0.04 0.07 0.07 0.08 0.09 0.07 0.07 0.09 0.07
RD5@1 0.75 0.41 0.46 0.67 0.4 0.43 0.38 0.3 0.48 0.54 0.39 0.59
RR5@1 0.02 0.07 0.05 0.09 0.07 0.05 0.05 0.05 0.06 0.09 0.07 0.11
@@
Table 15: Performance of code generation models against syntax perturbations. (HumanEval-X and worst case passes)
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS Nominal RP5@1 0.34 0.1 0.37 0.46 0.16 0.43 0.57 0.11 0.41 0.55 0.2 0.44 OperandSwap RP5@1 0.29 0.04 0.29 0.32 0.08 0.33 0.37 0.1 0.21 0.35 0.1 0.29 RD5@1 0.11 0.65 0.25 0.27 0.54 0.25 0.34 0.24 0.45 0.35 0.53 0.35 RR5@1 0.2 0.08 0.11 0.18 0.13 0.11 0.2 0.08 0.18 0.19 0.14 0.19 ForWhileTransformer RP5@1 0.26 0.04 0.23 0.3 0.05 0.27 0.37 0.1 0.22 0.32 0.13 0.25 RD5@1 0.22 0.65 0.41 0.33 0.71 0.39 0.34 0.19 0.44 0.4 0.42 0.43 RR5@1 0.21 0.08 0.17 0.21 0.15 0.17 0.23 0.09 0.2 0.21 0.12 0.23 VarRenamerNaive RP5@1 0.27 0.02 0.27 0.34 0.06 0.29 0.38 0.07 0.24 0.32 0.1 0.27 RD5@1 0.19 0.76 0.3 0.25 0.64 0.33 0.32 0.43 0.38 0.41 0.56 0.39 RR5@1 0.17 0.08 0.12 0.21 0.13 0.15 0.2 0.07 0.16 0.22 0.13 0.21 DeadCodeInserter RP5@1 0.04 0.01 0.08 0.05 0.02 0.1 0.07 0.02 0.06 0.07 0.04 0.09 RD5@1 0.89 0.88 0.79 0.89 0.89 0.78 0.87 0.81 0.84 0.88 0.81 0.81 RR5@1 0.29 0.09 0.3 0.4 0.15 0.34 0.49 0.12 0.33 0.47 0.19 0.38 VarRenamerRN RP5@1 0.26 0.03 0.26 0.31 0.04 0.3 0.38 0.07 0.24 0.31 0.09 0.26 RD5@1 0.22 0.71 0.32 0.3 0.75 0.32 0.33 0.48 0.39 0.42 0.61 0.42 RR5@1 0.2 0.07 0.12 0.22 0.15 0.14 0.22 0.09 0.15 0.23 0.13 0.22 VarRenamerCB RP5@1 0.27 0.03 0.27 0.33 0.07 0.32 0.39 0.05 0.22 0.37 0.1 0.26 RD5@1 0.17 0.71 0.29 0.26 0.57 0.28 0.3 0.57 0.44 0.32 0.56 0.42 RR5@1 0.21 0.09 0.11 0.2 0.15 0.12 0.21 0.07 0.17 0.18 0.13 0.23
@@
Table 16: Performance of code generation models against format perturbations.
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi
Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS
Nominal RP5@1 0.34 0.1 0.37 0.46 0.16 0.43 0.57 0.11 0.41 0.55 0.2 0.44
split_lines RP5@1 0.3 0.03 0.28 0.37 0.07 0.32 0.47 0.1 0.27 0.43 0.1 0.26
RD5@1 0.07 0.71 0.27 0.16 0.57 0.26 0.16 0.19 0.31 0.2 0.53 0.4
RR5@1 0.12 0.09 0.13 0.09 0.1 0.13 0.1 0.05 0.16 0.11 0.12 0.19
new_line_afterdoc RP5@1 0.25 0.04 0.2 0.38 0.09 0.37 0.5 0.12 0.3 0.46 0.14 0.21
RD5@1 0.24 0.65 0.48 0.14 0.5 0.17 0.11 0.1 0.22 0.14 0.36 0.51
RR5@1 0.12 0.07 0.18 0.11 0.1 0.07 0.09 0.02 0.12 0.1 0.08 0.25
new_line_aftercode RP5@1 0.2 0.04 0.18 0.36 0.07 0.35 0.42 0.09 0.16 0.45 0.12 0.18
RD5@1 0.39 0.59 0.52 0.19 0.61 0.19 0.25 0.29 0.59 0.16 0.47 0.58
RR5@1 0.16 0.07 0.21 0.16 0.12 0.11 0.16 0.04 0.26 0.11 0.1 0.29
tab_indent RP5@1 0.12 0.04 0.35 0.08 0.09 0.41 0.49 0.1 0.34 0.43 0.14 0.37
RD5@1 0.63 0.59 0.08 0.82 0.46 0.07 0.12 0.24 0.14 0.19 0.36 0.15
RR5@1 0.21 0.06 0.04 0.37 0.08 0.05 0.08 0.03 0.09 0.1 0.09 0.14
doc2comments RP5@1 0.1 0.03 0.39 0.2 0.06 0.02 0.35 0.1 0.38 0.4 0.12 0.41
RD5@1 0.69 0.71 -0.02 0.56 0.64 0.94 0.38 0.24 0.03 0.25 0.47 0.06
RR5@1 0.23 0.07 0.04 0.25 0.13 0.41 0.21 0.04 0.09 0.15 0.1 0.07
new_lines RP5@1 0.24 0.04 0.14 0.32 0.07 0.32 0.46 0.09 0.18 0.36 0.12 0.15
RD5@1 0.28 0.59 0.63 0.29 0.61 0.26 0.18 0.29 0.55 0.33 0.47 0.65
RR5@1 0.13 0.07 0.24 0.15 0.13 0.12 0.12 0.04 0.26 0.19 0.1 0.31

### v4

@@ 5.1 RQ1: How robust are LLMs across different programming languages?
Table 9: Performance summary of perturbations against code generation models using HumanEval-X test cases and worst case passes (5 out of 5).
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi Magicoder-7B QwenCode
Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS Java CPP JS Java CPP JS
Nominal RP5@1 0.02 0.1 0.08 0.11 0.12 0.13 0.13 0.12 0.13 0.16 0.14 0.16 0.65 0.57 0.53 0.57 0.59 0.62
Partial RP5@1 0.33 0.1 0.38 0.45 0.17 0.44 0.56 0.13 0.39 0.54 0.22 0.44 0.88 0.63 0.76 0.79 0.65 0.70
[…]
Table 10: Performance summary of perturbations using EvalPlus test cases and partial cases (3 out of 5)
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-multi CodeGen-6B-multi Magicoder7B QwenCode
Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS Java CPP JS Java CPP JS
Nominal RP5@1 0.02 0.08 0.07 0.11 0.1 0.12 0.14 0.12 0.11 0.15 0.13 0.14 0.63 0.55 0.51 0.55 0.57 0.6
Partial RP5@1 0.32 0.09 0.37 0.45 0.14 0.44 0.56 0.13 0.39 0.55 0.2 0.44 0.86 0.61 0.74 0.77 0.63 0.68
[…]
Table 11: Performance summary of perturbations using partial cases using HumanEval-X test cases and partial passes (3 out of 5).
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi Magicoder-7B QwenCode
Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS Java CPP JS Java CPP JS
Nominal RP5@1 0.02 0.1 0.08 0.11 0.12 0.13 0.13 0.12 0.13 0.16 0.14 0.16 0.65 0.57 0.53 0.57 0.59 0.62
Partial RP5@1 0.33 0.1 0.38 0.45 0.17 0.44 0.56 0.13 0.39 0.54 0.22 0.44 0.88 0.63 0.76 0.79 0.65 0.70
[…]
@@
Table 12: Performance of code generation models against docstring perturbations (worst case passes)
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-multi CodeGen-6B-multi Magicoder7B QwenCode
Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS Java CPP JS Java CPP JS
Nominal RP5@1 0.02 0.08 0.07 0.11 0.1 0.12 0.14 0.12 0.11 0.15 0.13 0.14 0.63 0.55 0.51 0.55 0.57 0.6
SynonymSubstitution RP5@1 0.01 0.01 0.02 0.02 0.04 0.02 0.05 0.04 0.02 0.04 0.06 0.05 0.36 0.33 0.26 0.16 0.31 0.31
RD5@1 0.5 0.85 0.75 0.78 0.62 0.79 0.65 0.65 0.78 0.72 0.52 0.65 0.43 0.4 0.49 0.7 0.45 0.48
RR5@1 0.02 0.07 0.05 0.11 0.06 0.09 0.09 0.08 0.09 0.11 0.07 0.1 0.29 0.23 0.27 0.41 0.27 0.32
ButterFingersPerturbation RP5@1 0.0 0.02 0.04 0.02 0.05 0.05 0.04 0.05 0.04 0.04 0.06 0.05 0.38 0.34 0.25 0.23 0.37 0.29
RD5@1 1.0 0.69 0.5 0.83 0.5 0.53 0.74 0.6 0.67 0.76 0.52 0.61 0.39 0.38 0.51 0.58 0.35 0.52
RR5@1 0.02 0.07 0.05 0.1 0.05 0.06 0.1 0.09 0.07 0.12 0.07 0.09 0.26 0.22 0.3 0.34 0.23 0.35
ChangeCharCase RP5@1 0.0 0.0 0.01 0.02 0.05 0.03 0.04 0.04 0.02 0.04 0.07 0.04 0.36 0.3 0.23 0.2 0.37 0.34
RD5@1 1.0 1.0 0.83 0.83 0.5 0.74 0.74 0.7 0.83 0.72 0.43 0.74 0.43 0.44 0.54 0.64 0.34 0.43
RR5@1 0.02 0.08 0.07 0.1 0.06 0.09 0.1 0.09 0.09 0.11 0.05 0.1 0.3 0.24 0.29 0.39 0.23 0.29
SynonymInsertion RP5@1 0.01 0.03 0.02 0.02 0.05 0.03 0.08 0.05 0.02 0.07 0.06 0.02 0.42 0.38 0.34 0.22 0.35 0.35
RD5@1 0.75 0.62 0.67 0.83 0.5 0.74 0.43 0.55 0.78 0.52 0.52 0.87 0.33 0.3 0.33 0.6 0.39 0.42
RR5@1 0.02 0.06 0.06 0.1 0.05 0.09 0.07 0.07 0.09 0.09 0.07 0.12 0.21 0.18 0.23 0.37 0.22 0.3
TenseTransformationFuture RP5@1 0.01 0.03 0.04 0.04 0.06 0.07 0.1 0.06 0.05 0.06 0.07 0.05 0.44 0.41 0.35 0.3 0.39 0.38
RD5@1 0.5 0.62 0.5 0.67 0.38 0.42 0.3 0.5 0.5 0.6 0.48 0.61 0.3 0.24 0.31 0.46 0.31 0.37
RR5@1 0.01 0.06 0.04 0.09 0.04 0.06 0.05 0.06 0.05 0.09 0.06 0.09 0.24 0.13 0.18 0.34 0.19 0.26
BackTranslation RP5@1 0.0 0.02 0.05 0.04 0.07 0.05 0.09 0.07 0.07 0.08 0.07 0.08 0.46 0.42 0.29 0.29 0.43 0.4
RD5@1 1.0 0.69 0.33 0.67 0.25 0.53 0.39 0.45 0.39 0.48 0.43 0.43 0.26 0.23 0.42 0.47 0.25 0.34
RR5@1 0.02 0.07 0.02 0.09 0.04 0.07 0.07 0.05 0.04 0.07 0.07 0.06 0.2 0.14 0.24 0.29 0.15 0.23
SwapCharactersPerturbation RP5@1 0.0 0.02 0.02 0.01 0.05 0.05 0.07 0.05 0.03 0.04 0.06 0.05 0.41 0.38 0.3 0.3 0.38 0.3
RD5@1 1.0 0.77 0.75 0.89 0.44 0.58 0.52 0.6 0.72 0.72 0.52 0.61 0.35 0.31 0.41 0.46 0.32 0.51
RR5@1 0.02 0.07 0.05 0.1 0.04 0.08 0.09 0.07 0.08 0.11 0.07 0.09 0.22 0.18 0.22 0.32 0.21 0.33
WhitespacePerturbation RP5@1 0.0 0.03 0.02 0.02 0.05 0.03 0.04 0.05 0.03 0.04 0.05 0.04 0.37 0.32 0.25 0.27 0.33 0.37
RD5@1 1.0 0.62 0.75 0.78 0.5 0.74 0.7 0.6 0.72 0.76 0.57 0.7 0.41 0.42 0.51 0.5 0.42 0.39
RR5@1 0.02 0.07 0.05 0.1 0.05 0.1 0.1 0.07 0.08 0.12 0.07 0.1 0.27 0.23 0.27 0.34 0.25 0.27
TenseTransformationPast RP5@1 0.0 0.04 0.04 0.04 0.05 0.06 0.08 0.07 0.07 0.04 0.07 0.06 0.51 0.43 0.37 0.32 0.43 0.38
RD5@1 1.0 0.54 0.42 0.67 0.5 0.47 0.43 0.45 0.39 0.72 0.48 0.57 0.19 0.21 0.27 0.41 0.25 0.37
RR5@1 0.02 0.05 0.03 0.09 0.05 0.05 0.06 0.07 0.04 0.11 0.06 0.08 0.16 0.14 0.16 0.27 0.16 0.25
EnglishInflectionalVariation RP5@1 0.0 0.02 0.04 0.02 0.04 0.06 0.04 0.05 0.04 0.06 0.05 0.06 0.48 0.41 0.34 0.35 0.4 0.37
RD5@1 1.0 0.69 0.5 0.83 0.56 0.47 0.7 0.55 0.61 0.6 0.57 0.57 0.24 0.24 0.34 0.36 0.3 0.39
RR5@1 0.02 0.07 0.04 0.09 0.05 0.05 0.1 0.07 0.07 0.1 0.07 0.08 0.21 0.17 0.2 0.27 0.17 0.25
@@
Table 13: Performance of code generation models against function name perturbations. (worst case passes)
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-multi CodeGen-6B-multi Magicoder7B QwenCode
Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS Java CPP JS Java CPP JS
Nominal RP5@1 0.02 0.08 0.07 0.11 0.1 0.12 0.14 0.12 0.11 0.15 0.13 0.14 0.63 0.55 0.51 0.55 0.57 0.6
FuncRenameSwapChar RP5@1 0.01 0.04 0.05 0.05 0.06 0.05 0.09 0.05 0.06 0.05 0.09 0.08 0.52 0.41 0.39 0.3 0.47 0.39
RD5@1 0.75 0.54 0.25 0.56 0.38 0.53 0.39 0.55 0.44 0.64 0.33 0.43 0.17 0.24 0.23 0.46 0.17 0.35
RR5@1 0.02 0.05 0.03 0.07 0.04 0.06 0.05 0.08 0.05 0.1 0.05 0.09 0.11 0.15 0.16 0.29 0.1 0.23
FuncRenameChangeChar RP5@1 0.0 0.01 0.07 0.04 0.06 0.07 0.07 0.03 0.07 0.05 0.06 0.08 0.53 0.44 0.41 0.29 0.49 0.4
RD5@1 1.0 0.85 0.08 0.61 0.38 0.42 0.52 0.75 0.39 0.68 0.52 0.43 0.16 0.2 0.18 0.47 0.13 0.34
RR5@1 0.02 0.09 0.03 0.08 0.04 0.06 0.07 0.09 0.05 0.12 0.08 0.07 0.12 0.16 0.15 0.27 0.12 0.22
FuncRenameSynonymSub RP5@1 0.0 0.02 0.04 0.04 0.05 0.07 0.08 0.06 0.04 0.07 0.07 0.08 0.48 0.42 0.37 0.3 0.46 0.42
RD5@1 1.0 0.69 0.42 0.67 0.44 0.42 0.43 0.5 0.67 0.56 0.48 0.43 0.24 0.23 0.28 0.46 0.19 0.3
RR5@1 0.02 0.07 0.04 0.07 0.04 0.05 0.06 0.07 0.07 0.09 0.06 0.07 0.18 0.14 0.18 0.3 0.13 0.23
FuncRenameInflectionalVariation RP5@1 0.0 0.01 0.05 0.04 0.06 0.06 0.1 0.05 0.07 0.04 0.07 0.08 0.49 0.45 0.37 0.32 0.45 0.41
RD5@1 1.0 0.85 0.33 0.61 0.38 0.47 0.26 0.6 0.39 0.72 0.43 0.43 0.22 0.19 0.27 0.42 0.2 0.32
RR5@1 0.02 0.07 0.02 0.07 0.05 0.07 0.05 0.07 0.05 0.11 0.05 0.06 0.14 0.13 0.17 0.27 0.12 0.22
FuncRenameButterFinger RP5@1 0.0 0.02 0.04 0.02 0.04 0.06 0.09 0.07 0.07 0.04 0.07 0.07 0.51 0.47 0.37 0.29 0.43 0.41
RD5@1 1.0 0.69 0.42 0.78 0.62 0.47 0.39 0.4 0.39 0.72 0.48 0.48 0.18 0.14 0.27 0.47 0.25 0.32
RR5@1 0.02 0.08 0.03 0.09 0.06 0.07 0.05 0.05 0.05 0.11 0.06 0.08 0.12 0.13 0.2 0.3 0.15 0.23
FuncRenameSnake/CamelCase RP5@1 0.01 0.04 0.05 0.04 0.07 0.07 0.09 0.07 0.06 0.07 0.07 0.08 0.53 0.45 0.39 0.32 0.44 0.4
RD5@1 0.75 0.54 0.25 0.61 0.31 0.42 0.39 0.4 0.44 0.52 0.43 0.43 0.16 0.18 0.23 0.41 0.23 0.33
RR5@1 0.02 0.07 0.03 0.08 0.05 0.05 0.05 0.05 0.05 0.08 0.05 0.07 0.13 0.11 0.15 0.3 0.13 0.21
Table 14: Performance of code generation models against syntax perturbations. (worst case passes)
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-multi CodeGen-6B-multi Magicoder7B QwenCode
Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS Java CPP JS Java CPP JS
Nominal (Partial) RP5@1 0.32 0.09 0.37 0.45 0.14 0.44 0.56 0.13 0.39 0.55 0.2 0.44 0.86 0.61 0.74 0.77 0.63 0.68
VarRenamerRN RP5@1 0.25 0.01 0.27 0.28 0.03 0.3 0.36 0.04 0.26 0.29 0.06 0.27 0.68 0.48 0.6 0.66 0.41 0.41
RD5@1 0.23 0.86 0.28 0.37 0.78 0.31 0.36 0.67 0.34 0.47 0.69 0.39 0.21 0.21 0.19 0.14 0.35 0.39
RR5@1 0.2 0.07 0.12 0.24 0.13 0.13 0.21 0.1 0.13 0.26 0.13 0.21 0.21 0.18 0.19 0.39 0.3 0.33
VarRenamerNaive RP5@1 0.27 0.02 0.27 0.32 0.05 0.28 0.37 0.06 0.25 0.3 0.07 0.28 0.7 0.53 0.62 0.63 0.46 0.52
RD5@1 0.17 0.79 0.26 0.29 0.61 0.36 0.35 0.52 0.36 0.44 0.62 0.36 0.19 0.13 0.17 0.17 0.28 0.23
RR5@1 0.18 0.08 0.11 0.21 0.11 0.16 0.21 0.09 0.15 0.24 0.12 0.2 0.21 0.14 0.19 0.35 0.25 0.29
ForWhileTransformer RP5@1 0.24 0.02 0.23 0.27 0.04 0.28 0.38 0.07 0.24 0.33 0.1 0.26 0.68 0.52 0.56 0.68 0.42 0.39
RD5@1 0.25 0.71 0.38 0.38 0.7 0.36 0.33 0.43 0.39 0.4 0.47 0.42 0.21 0.15 0.25 0.11 0.34 0.43
RR5@1 0.21 0.09 0.16 0.22 0.12 0.16 0.2 0.09 0.18 0.22 0.1 0.22 0.2 0.14 0.23 0.29 0.29 0.34
OperandSwap RP5@1 0.28 0.02 0.27 0.32 0.07 0.33 0.38 0.07 0.23 0.34 0.09 0.28 0.64 0.54 0.65 0.62 0.49 0.5
RD5@1 0.13 0.71 0.26 0.27 0.52 0.25 0.33 0.43 0.41 0.38 0.53 0.36 0.26 0.11 0.12 0.19 0.22 0.27
RR5@1 0.2 0.09 0.11 0.17 0.11 0.11 0.2 0.09 0.16 0.21 0.13 0.2 0.22 0.15 0.14 0.37 0.21 0.27
DeadCodeInserter RP5@1 0.04 0.01 0.09 0.05 0.01 0.1 0.07 0.01 0.07 0.07 0.02 0.1 0.04 0.13 0.09 0.03 0.12 0.07
RD5@1 0.89 0.93 0.75 0.89 0.91 0.76 0.87 0.95 0.83 0.88 0.88 0.78 0.96 0.78 0.88 0.96 0.81 0.9
RR5@1 0.3 0.08 0.28 0.4 0.13 0.34 0.49 0.12 0.32 0.48 0.17 0.37 0.82 0.48 0.66 0.74 0.51 0.62
VarRenamerCB RP5@1 0.26 0.02 0.28 0.3 0.06 0.32 0.36 0.04 0.23 0.35 0.07 0.26 0.66 0.48 0.65 0.65 0.42 0.44
RD5@1 0.19 0.79 0.25 0.33 0.57 0.26 0.36 0.71 0.42 0.36 0.66 0.4 0.23 0.21 0.13 0.16 0.34 0.36
RR5@1 0.22 0.08 0.1 0.22 0.13 0.12 0.21 0.09 0.16 0.21 0.13 0.23 0.22 0.2 0.16 0.4 0.29 0.28
Table 15: Performance of code generation models against format perturbations (worst case passes).
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-multi CodeGen-6B-multi Magicoder7B QwenCode
Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS Java CPP JS Java CPP JS
Nominal (Partial) RP5@1 0.32 0.09 0.37 0.45 0.14 0.44 0.56 0.13 0.39 0.55 0.2 0.44 0.86 0.61 0.74 0.77 0.63 0.68
split_lines RP5@1 0.01 0.02 0.26 0.39 0.06 0.33 0.02 0.07 0.27 0.03 0.09 0.27 0.77 0.49 0.59 0.7 0.49 0.49
RD5@1 0.98 0.71 0.3 0.12 0.57 0.25 0.97 0.43 0.31 0.94 0.56 0.37 0.1 0.19 0.21 0.09 0.22 0.29
RR5@1 0.32 0.07 0.13 0.07 0.08 0.12 0.54 0.07 0.15 0.52 0.11 0.18 0.11 0.14 0.21 0.15 0.18 0.26
doc2comments RP5@1 0.11 0.01 0.36 0.2 0.05 0.02 0.34 0.09 0.36 0.38 0.1 0.4 0.79 0.52 0.72 0.48 0.52 0.73
RD5@1 0.66 0.86 0.03 0.55 0.65 0.94 0.4 0.33 0.08 0.3 0.47 0.1 0.08 0.14 0.03 0.37 0.17 -0.06
RR5@1 0.23 0.07 0.05 0.24 0.12 0.41 0.23 0.05 0.08 0.18 0.1 0.08 0.09 0.13 0.11 0.34 0.11 0.2
new_lines RP5@1 0.23 0.02 0.15 0.3 0.06 0.34 0.45 0.07 0.23 0.35 0.1 0.18 0.77 0.52 0.68 0.8 0.5 0.51
RD5@1 0.3 0.71 0.61 0.32 0.57 0.22 0.2 0.43 0.41 0.36 0.47 0.6 0.11 0.15 0.09 -0.04 0.21 0.26
RR5@1 0.13 0.07 0.24 0.16 0.1 0.11 0.12 0.07 0.21 0.2 0.1 0.29 0.1 0.13 0.1 0.18 0.16 0.23
new_line_afterdoc RP5@1 0.26 0.02 0.2 0.39 0.06 0.38 0.51 0.09 0.32 0.45 0.12 0.23 0.81 0.51 0.71 0.68 0.52 0.57
RD5@1 0.21 0.71 0.46 0.12 0.57 0.14 0.1 0.29 0.17 0.18 0.38 0.47 0.06 0.16 0.05 0.12 0.17 0.17
RR5@1 0.1 0.06 0.18 0.1 0.09 0.06 0.08 0.04 0.1 0.11 0.07 0.23 0.06 0.11 0.07 0.15 0.11 0.18
new_line_aftercode RP5@1 0.19 0.02 0.18 0.35 0.05 0.34 0.41 0.07 0.2 0.42 0.1 0.2 0.76 0.52 0.62 0.77 0.54 0.45
RD5@1 0.42 0.71 0.51 0.21 0.61 0.22 0.26 0.43 0.5 0.23 0.5 0.54 0.12 0.14 0.16 -0.01 0.14 0.34
RR5@1 0.17 0.07 0.2 0.14 0.09 0.11 0.16 0.05 0.23 0.14 0.11 0.27 0.12 0.13 0.15 0.2 0.15 0.26
tab_indent RP5@1 0.12 0.03 0.34 0.07 0.08 0.41 0.48 0.08 0.32 0.43 0.12 0.38 0.81 0.48 0.71 0.73 0.49 0.57
RD5@1 0.64 0.64 0.08 0.84 0.43 0.07 0.15 0.38 0.19 0.22 0.41 0.12 0.06 0.21 0.05 0.05 0.22 0.17
RR5@1 0.21 0.05 0.05 0.37 0.06 0.05 0.1 0.06 0.11 0.12 0.09 0.13 0.07 0.16 0.07 0.24 0.15 0.25